In a triangle, one angle is 6 degrees more than the second angle. the third angle is 18 degrees more than the second angle. Find the measure of each angle.
Answer by mananth(16949)

let the second angle be x
first angle will be x+6
Third angle will be 18+(x+6)= x+24
x+x+6+x+24=180
3x+30=180
3x=150
x= 50 second angle
first angle = 56
third angle = 50+24= 74 deg